# CLIP-EBC: CLIP Can Count Accurately through Enhanced Blockwise Classification

### v2

@@ Abstract
We propose CLIP-EBC, the first fully CLIP-based model for accurate crowd density estimation. While the CLIP model has demonstrated remarkable success in addressing recognition tasks such as zero-shot image classification, its potential for counting has been largely unexplored due to the inherent challenges in transforming a regression problem, such as counting, into a recognition task. In this work, we investigate and enhance CLIP’s ability to count, focusing specifically on the task of estimating crowd sizes from images. Existing classification-based crowd-counting frameworks have significant limitations, including the quantization of count values into bordering real-valued bins and the sole focus on classification errors. These practices result in label ambiguity near the shared borders and inaccurate prediction of count values. Hence, directly applying CLIP within these frameworks may yield suboptimal performance. To address these challenges, we first propose the Enhanced Blockwise Classification (EBC) framework. Unlike previous methods, EBC utilizes integer-valued bins, effectively reducing ambiguity near bin boundaries. Additionally, it incorporates a regression loss based on density maps to improve the prediction of count values. Within our backbone-agnostic EBC framework, we then introduce CLIP-EBC to fully leverage CLIP’s recognition capabilities for this task. Extensive experiments demonstrate the effectiveness of EBC and the competitive performance of CLIP-EBC. Specifically, our EBC framework can improve existing classification-based methods by up to 44.5% on the UCF-QNRF dataset, and CLIP-EBC achieves state-of-the-art performance on the NWPU-Crowd test set, with an MAE of 58.2 and an RMSE of 268.5, representing improvements of 8.6% and 13.3% over the previous best method, STEERER. The code and weights will be made publicly available.
1 Introduction
Figure 1: Comparison of our model CLIP-EBC (bottom right) with other CLIP-based crowd counting models: CrowdCLIP (top right) (Liang et al. 2023) and CLIP-Count (bottom left) (Jiang, Liu, and Chen 2023). The ground-truth count for the input image (top left) is 177. CLIP-EBC accurately predicts 178.1, while CrowdCLIP and CLIP-Count significantly overestimate with predictions of 780.0 and 673.2, respectively. CLIP-EBC is also unique in being purely CLIP-based and capable of generating density maps, unlike CrowdCLIP and CLIP-Count, which either estimates the global count value or incorporates additional modules.
Crowd counting aims to automatically estimate the number of people present in an image or a video. In recent years, this task has garnered considerable attention due to its potential applications in critical areas such as epidemic control (Valencia et al. 2021) and crowd management (Saleh, Suandi, and Ibrahim 2015). Several crowd datasets (Zhang et al. 2016; Idrees et al. 2018; Sindagi, Yasarla, and Patel 2020; Wang et al. 2020b) have been proposed in response to the growing interest. These datasets typically provide crowd images annotated with sparse binary density maps, where a value of 1 indicates a labeled head center, and 0 indicates otherwise. The spatial size of each person in the image is not labeled due to the heavy workload and occlusion. Therefore, many contemporary approaches (Li, Zhang, and Chen 2018; Liu, Salzmann, and Fua 2019; Song et al. 2021b; Ma, Sanchez, and Guha 2022) adopt an encoder-decoder framework, aiming to directly regress the density maps. However, the sparsity in the ground-truth density maps makes training difficult. To address this challenge, these methods usually employ 1) Gaussian smoothing as preprocessing; and 2) blockwise prediction by generating density maps of reduced spatial sizes, where each value represents the estimated count in the corresponding block of the image.
@@ 1 Introduction
Figure 2: Overview of our CLIP-EBC model implemented using our proposed EBC framework with fine-grained bins. Given an input image, we utilize the CLIP (Radford et al. 2021) image encoder to extract feature maps. At top right of this figure, the predefined bins (e.g. $\{0\}$) and the template (e.g. “There is {CLS} person”) are used to construct text prompts. We freeze the CLIP text encoder and employ it to extract text embeddings from the prompts. For the image feature vector at each spatial location, we calculate its cosine similarity with the text embedding for each bin. The probability map is generated by applying softmax to these similarities, and the density map is obtained by averaging the preset values (e.g. 1 for bin $\{1\}$) weighted by the probability scores. To fine-tune the image encoder, we leverage the proposed $\mathcal{L}_{\text{DACE}}$ loss (Eq. (3.1)).
Although CLIP (Contrastive Language-Image Pretraining) (Valencia et al. 2021) has achieved significant success in recognition-related downstream tasks such as object detection (Gu et al. 2022) and semantic segmentation (Lüddecke and Ecker 2022), its application to crowd counting remains largely unexplored. This gap arises mainly from the inherent mismatch between CLIP, which is originally designed for recognition, and counting, which constitutes a regression task. While there have been some attempts (Liu et al. 2019; Xiong et al. 2019) to reframe crowd counting as a blockwise classification problem, these approaches tend to exhibit several notable limitations. For instance, similar to their regression-based counterparts, these methods employ Gaussian smoothing to preprocess the ground-truth density maps, transforming the initially discrete count value space $\mathbb{N}$ into a continuous space $\mathbb{R}_{+}$. Consequently, these methods use a sequence of bordering real-valued intervals as bins to cover this space, but the lack of separation between adjacent bins leads to significant label ambiguity, hindering the classification of samples near bin boundaries. For example, distinguishing whether a block with a local count value of 0.05 belongs in the bin $[0,0.05)$ or $[0.05,0.1)$ becomes challenging. Also, Gaussian smoothing introduces noise into labels since head sizes are not annotated (Wang et al. 2020a). Another limitation of current classification-based methods is their sole focus on the classification error without considering the proximity of predicted count values, i.e., the expectation over the bins, to the ground truth. Since low classification error does not necessarily correlate with low count error, their loss functions may lead to inaccurate count estimation.
To address these challenges, we propose an Enhanced Blockwise Classification (EBC) framework, specifically designed to overcome the limitations of current classification-based methods. Our EBC framework is model-agnostic in the sense that existing regression-based methods can be seamlessly integrated into the EBC framework with minimal modifications, resulting in significant performance enhancements. Building upon the EBC framework, we explore the potential of leveraging the original structure of CLIP for crowd counting and introduce CLIP-EBC. Compared to other CLIP-based methods (Liang et al. 2023; Jiang, Liu, and Chen 2023) (see Fig. 1), CLIP-EBC (detailed in Fig. 2) stands out as the first fully CLIP-based model capable of accurately estimating crowd sizes. Experimental results across four datasets demonstrate the substantial improvement offered by our EBC framework over existing regression-based methods. EBC exhibits notable efficacy, achieving up to a 44.5% reduction in Mean Absolute Error (MAE) for the standard blockwise classification method (Liu et al. 2019) on the UCF-QNRF (Idrees et al. 2018) dataset. Furthermore, our proposed CLIP-EBC model surpasses state-of-the-art crowd-counting methods, showcasing its effectiveness. Specifically, CLIP-EBC achieves an MAE of 58.2 and a RMSE of 268.5 on the NWPU-Crowd (Wang et al. 2020b) test split. These results demonstrate that CLIP, with the support of the EBC framework, can accurately estimate crowd density distributions. To summarize, our contributions are as follows:
@@
  • We propose the novel Enhanced Blockwise Classifica-tion (EBC) framework that substantially improves over the classification-based methods by addressing issues related to Gaussian smoothing and loss functions.
@@
  • Building upon EBC, we present the first fully CLIP-based crowd counting model CLIP-EBC that is caple of accurately estimating crowd distributions.
@@
2 Related Work
2.1 Regression-Based Methods
Crowd counting is dominated by a large variety of encoder-decoder-based models that regress density maps. Some models focus on tackling the scale variation problem caused by perspective distortion. Zhang et al.(Zhang et al. 2016) introduce a multi-column CNN structure. Each branch extracts feature maps of different receptive field sizes, which are subsequently fused through concatenation. In contrast, Liu et al.(Liu, Salzmann, and Fua 2019) propose a VGG-16 (Simonyan and Zisserman 2015)-based single-branch model, incorporating a multi-scale module to extract and fuse features across scales. Recognizing the importance of enlarging receptive fields, Li et al.(Li, Zhang, and Chen 2018) advocate for the use of dilated convolutions in generating density maps. Since ground-truth density maps are often sparse, smoothing them with Gaussian kernels is a common strategy to facilitate model optimization. Nevertheless, this approach introduces the challenge of selecting appropriate kernel widths, which ideally should match scales. Unfortunately, head sizes are usually not provided for crowd-counting tasks. Hence, Gaussian smoothing unavoidably introduces errors and noise in labels: if the kernel size is set too small, those pixels that correspond to individual’s heads are set to 00 in the density maps; conversely, if the kernel size is too large, those pixels that correspond to the background can be mistaken as pedestrians. To address this issue, Wang et al.(Wang et al. 2020a) introduce the DMCount loss by leveraging the discrete optimal transport theory. This loss function does not require Gaussian smoothing and has been shown to enhance performance.
2.2 Classification-Based Methods
Existing classification-based approaches emerge from a motivation to rectify the long-tail distribution of count values, where large values are gravely underrepresented, adversely affecting the performance of regression-based models. To address this issue, classification-based methods partition the support range $[0,\,\infty)$ into non-overlapping intervals to enhance the sample size for each class. During inference, the middle points of each interval, weighted by the probability score, are added up as the predicted count. For example, Xiong et al.(Xiong et al. 2019) introduce DCNet, which predicts counts at multiple levels by using the same set of bins. However, this method neglects the fact that large values are less likely to appear at a local level, thereby exacerbating class imbalance. To handle this issue, Liu et al.(Liu et al. 2019) propose the concept of blockwise classification with models outputting probability maps, where the vector at each pixel represents the predicted probability scores. However, similar to their regression-based counterparts, these methods also smooth the ground-truth density maps with Gaussian kernels, introducing noise in the labels. Another problem with Gaussian smoothing is that it maps count values from a discrete integer space into a continuous real space, which can be partitioned only by intervals (or bins) with shared borders, e.g., $(0,0.5]$ and $(0.5,1]$. This quantization policy makes it difficult to classify the sample points near the boundaries of the intervals. Moreover, these methods solely focus on classification results, overlooking the fact that two probability distributions can have the same classification error but different expectations, thus severely impacting the performance during testing.
2.3 CLIP in Crowd Counting
The Contrastive Language-Image Pre-training (CLIP) model (Radford et al. 2021) has demonstrated outstanding performance in downstream tasks such as zero-shot image classification (Radford et al. 2021), object detection (Gu et al. 2022) and (Lüddecke and Ecker 2022), but there are only few studies in crowd counting. Liang et al.(Liang et al. 2023) propose an unsupervised method established on ranking. Their method utilizes sequences of nested image patches and a predefined sequence of count numbers as input, with the objective of minimizing the ranking loss of similarities. Nevertheless, one limitation of this approach is its inability to generate density maps, crucial for applications such as pandemic control and public safety monitoring. Jiang et al.(Jiang, Liu, and Chen 2023) introduce a text-guided zero-shot counting method, which leverages CLIP to generate text guidance, yet it remains rooted in density-map regression, leading to suboptimal performance in crowd counting tasks. In contrast, we are the first to investigate the potential of using only CLIP to estimate crowd density maps. To this end, we introduce CLIP-EBC, the first crowd-counting method fully based on CLIP, and demonstrate that CLIP can accurately estimate crowd distributions while retaining its structure as much as possible.
3 Methods
@@
Our Enhanced Blockwise Classification (EBC) framework improves the Standard Blockwise Classification (SBC) (Liu et al. 2019) from three key aspects: the bin strategy, noise control, and loss formulation. In this section, we first briefly recapitulate SBC, and then provide a detailed explanation of our enhancements. Finally, we demonstrate how to leverage our backbone-agnostic EBC framework with CLIP for density map estimation, introducing our model CLIP-EBC.
3.1 Enhanced Blockwise Classification
Recap: Standard Blockwise Classification (SBC).
@@ 3.1 Enhanced Blockwise Classification
Similar to many regression-based methods (Li, Zhang, and Chen 2018; Xiong et al. 2019; Liu et al. 2019; Ma et al. 2019; Wang et al. 2020a; Ma, Sanchez, and Guha 2022), SBC (Liu et al. 2019) employs blockwise prediction. Formally, given an input image $\boldsymbol{X}\in\mathbb{R}_{+}^{3\times H\times W}$, with $3$ channels of height $H$ and width $W$, SBC outputs a probability map
@@
where ///// / represents the floor division operator and the integer $r$ is a backbone-related reduction factor. At each spatial location $(i,\,j)$, the $n$-dimensional vector $\boldsymbol{P}^{*}_{:,\,i,\,j}$ represents the predicted probability distribution of the count value in the region $(r(i-1):ri,\,r(j-1):rj)$ of the image over the $n$ predefined bins $\{\mathcal{B}_{i}\mid i=1,\cdots,n\}$. These bins satisfy
Condition 1: $\quad\forall p\neq q,\,\mathcal{B}_{p}\cap\mathcal{B}_{q}=\emptyset;$
Condition 2: $\quad\mathcal{S}\subseteq\cup_{i=1}^{n}\mathcal{B}_{i},$
@@
Enhancement 1: Integer-Valued Bins.
@@
(a) Original area around the box
(b) Zoomed-in view of the area
Figure 3: An example from ShanghaiTech A dataset (Zhang et al. 2016) illustrating an extremely dense area with erroneous annotations. The $8\times 8$ magenta box in Fig. 3(a) highlights a congested region with an labeled count of 9 individuals. However, the zoomed-in view in Fig. 3(b) reveals that no discernible human figures can be identified within the marked box, evidencing label noise in highly dense regions.
Enhancement 2: Noise Reduction.
@@
Enhancement 3: Combined Loss.
Previous blockwise classification methods (Liu et al. 2019; Xiong et al. 2019) primarily focus on minimizing the cross-entropy loss between the predicted probability map $\boldsymbol{P}^{*}$ and the ground truth $\boldsymbol{P}$ but often overlook the mismatch in count values. Since two different probability distributions can yield the same classification error while having different expectations (or modes) for the predicted count values, models trained with such loss functions may not guarantee accurate count estimates. To address this limitation, we propose the Distance-Aware-Cross-Entropy (DACE) loss:
@@
where $\boldsymbol{1}$ is the indicator function, $\boldsymbol{P}$ is the one-hot encoded ground-truth probability map, $\boldsymbol{P}^{*}$ is the predicted probability map, $\boldsymbol{Y}$ is the ground-truth density map, $\boldsymbol{Y}^{*}$ is the predicted density map, and $\lambda$ is a weighting factor balancing $\mathcal{L}_{\text{class}}$ and $\mathcal{L}_{\text{count}}$. Since the ground-truth density maps are not smoothed, we use the DMCount (Wang et al. 2020a) based on the discrete Wasserstein distance as the count loss $\mathcal{L}_{\text{count}}$.
Method E1 E2 E3 MAE $\downarrow$
SBC 140.6
+E1 88.3
+E1+E2 85.8
EBCa $\lambda=0.01$ 83.9
EBCb $\lambda=0.10$ 83.4
EBCc $\lambda=1.00$ 77.9
EBCd $\lambda=2.00$ 82.3
Table 1: Influence of each enhancement strategy in our EBC framework evaluated on UCF-QNRF (Idrees et al. 2018), where SBC represents the standard blockwise classification framework (Liu et al. 2019). E1, E2, and E3 denote our three enhancement strategies: integer-valued bins, noise reduction and combined loss. Note that EBC=SBC+E1+E2+E3. The reduced mean abolute errors demonstrate the effectivness of our enhancement strategies and the superiority of EBC.
3.2 The Structure of CLIP-EBC
Fig. 2 depicts the structure of our CLIP-EBC model with fine-grained bins (see Eq. (3)). As the mechanism of CLIP-EBC is based on our proposed EBC framework, here we focus on how to generate the predicted probability map $\boldsymbol{P}^{*}$. Additional essential details, such as the loss function used in training, are provided in Section 3.1.
where $d$ is the dimension of the CLIP embedding space.
@@ 3.2 The Structure of CLIP-EBC
Table 2: Our EBC framework can improve both classifica-tion-based and regression-based methods. We reimplement both SBC models (Liu et al. 2019) and two regression-based methods (CSRNet (Li, Zhang, and Chen 2018) and DMCount (Wang et al. 2020a)) using our EBC framework on the UCF-QNRF dataset. Regardless of the choice of the backbone, our EBC always improves the performance.

### v3

@@ Abstract
We propose CLIP-EBC, the first fully CLIP-based model designed for accurate crowd density estimation. Existing classification-based crowd counting frameworks face challenges when directly applied with CLIP. For instance, these methods quantize count values into bordering real-valued bins, which are inconsistent with CLIP’s pretraining corpus. Besides, this quantization strategy also introduces label ambiguity near shared bin borders. To address these issues, we propose the Enhanced Blockwise Classification (EBC) framework, which utilizes integer-valued bins and additionally incorporates a density-map-based loss to further improve count accuracy. Building on this backbone-agnostic framework, CLIP-EBC fully leverages CLIP’s recognition capabilities for crowd counting. Experiments show that EBC improves classification-based methods by up to 44.5% in mean absolute error (MAE) on the UCF-QNRF dataset. Furthermore, CLIP-EBC achieves state-of-the-art performance on the NWPU-Crowd test set, with an MAE of 58.2, surpassing the previous best method by 8.6%. Our code is available at https://github.com/Yiming-M/CLIP-EBC.
Index Terms:
crowd counting, crowd density estimation, CLIP
I Introduction
The vision-language model CLIP [1] is pre-trained on millions of image-text pairs using contrastive learning, enabling it to align visual and textual modalities effectively. This capability has allowed CLIP to excel in pattern recognition, setting new benchmarks in downstream tasks such as zero-shot image classification [1], and semantic segmentation [2]. Despite its strong performance in these recognition-based tasks, its potential for regression tasks, which require predicting continuous values, remains largely unexplored. This limitation arises primarily from its contrastive learning objective, which is inherently designed for discrete decision-making tasks, such as classification or retrieval, where the goal is to match or rank embeddings. However, regression tasks require predicting continuous values, which is not the focus of contrastive learning.
In this work, we aim to address this gap by investigating whether CLIP can effectively perform regression tasks, with a specific focus on crowd counting. Crowd counting is a particularly demanding task due to the inherent challenges, such as high-density crowds, occlusions, and scale variations. Thus, it provides a robust testbed to evaluate CLIP’s applicability to regression problems.
Figure 1: Comparison of our model CLIP-EBC and other CLIP-based crowd counting methods. Top: CrowdCLIP [3] predicts total counts from global image features. Due to the incapability to capture spatial details, it suffers from reduced accuracy. Middle: CLIP-Count [4] employs extra modules to generate density maps, improving accuracy at the cost of increased complexity. Bottom: CLIP-EBC is the first fully CLIP-based density-map approach, achieving high accuracy without additional modules.
Crowd counting is typically formulated as a blockwise density map regression task [5, 6, 7]. Therefore, an intuitive approach to adapt CLIP for crowd counting is through classification frameworks, where blockwise count values are quantized into discrete bins. However, directly applying CLIP in existing classification frameworks [8, 9] presents significant challenges.
@@ I Introduction
Figure 2: Overview of the EBC framework and the CLIP-EBC model. The EBC framework introduces an integer-based quantization strategy and noise reduction to partition count values into $n$ predefined bins (e.g., $\{0\},\{1\},\cdots,[m,\infty)$). Given an input image, the crowd counting model implemented under EBC generates a blockwise probability map. The density map is obtained by averaging the bin midpoints weighted by the probability map. Finally, the predicted probability map and the density map are fed into the proposed $\mathcal{L}_{\text{DACE}}$ loss (Eq. (III-A4)), which is used to fine-tune the model. CLIP-EBC, an instance of EBC, leverages the predefined bins and the prompt template to construct $n$ text prompts (e.g., “There is 0 person”). The frozen CLIP [1] text encoder extracts text embeddings, while the fine-tuned CLIP image encoder generates image feature maps. For the image feature vector at each spatial location, we calculate its cosine similarity with the text embedding for each bin and then apply softmax to obtain the probability map. Subsequent steps align with the EBC framework to generate the final outputs.
For instance, existing classification methods employ Gaussian smoothing to reduce the sparsity of the ground-truth density maps, transforming the initially discrete count value spaces $\mathbb{N}$ into a continuous space $\mathbb{R}_{+}$. This requires dividing the space into consecutive real-valued intervals to construct classification bins. While such bins can cover $\mathbb{R}_{+}$, they are poorly suited for constructing CLIP-compatible text prompts, since CLIP’s text encoder is pre-trained on natural language corpora, where descriptions are typically human-readable and discrete (e.g., “there are 3 people”). Prompts involving real values (e.g., “there are 3.14 people”) lack semantic clarity and may result in noisy text embeddings and degraded performance. Besides, the use of adjacent bins introduces label ambiguity, particularly for samples near shared boundaries. For example, the bins [0, 0.05) and [0.05, 0.1) in [8] create uncertainty for a block with a count of 0.05, as it is unclear which bin it should belong to. Such ambiguity weakens the reliability of predictions. Another limitation of current classification-based methods is their sole focus on optimizing the classification accuracy. This can lead to suboptimal performance on regression metrics, such as Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE), which are critical for counting tasks.
To address these challenges, we propose the Enhanced Blockwise Classification (EBC) framework, specifically designed to overcome the limitations of current classification-based methods. Our EBC framework is model-agnostic: existing regression-based methods can be implemented using EBC with minimal modifications to achieve substantial performance improvements. Building on this framework, we introduce CLIP-EBC, which fully leverages the original structure of CLIP for crowd counting. Unlike other CLIP-based methods [3, 4] (see Fig. 1), CLIP-EBC is the first model to fully utilize CLIP for density-map-based crowd counting. While CrowdCLIP [3] also relies solely on CLIP, it predicts total counts instead of density maps, limiting its ability to capture fine-grained spatial details for accurate predictions.
Experiments across four benchmark datasets demonstrate the significant advantages of our EBC framework. For instance, EBC reduces Mean Absolute Error (MAE) by up to 44.5% on the UCF-QNRF dataset [10] compared to the standard blockwise classification method [8]. Furthermore, CLIP-EBC surpasses state-of-the-art crowd counting methods, achieving an MAE of 58.2 and a Root Mean Square Error (RMSE) of 268.5 on the NWPU-Crowd test split [11]. These results demonstrate that CLIP, with the support of the EBC framework, can accurately estimate crowd density distributions. To summarize, our contributions are as follows:
@@
  • We propose the novel Enhanced Blockwise Classification (EBC) framework that substantially improves over the classification-based methods by addressing issues related to Gaussian smoothing and loss functions.
@@
  • Building upon EBC, we present the first fully CLIP-based crowd counting model CLIP-EBC that is capable of accurately estimating crowd densities.
@@
II Related Work
Classification-Based Crowd Counting Methods. These models aim to address the long-tail distribution of count values, where large counts are underrepresented, hindering the performance of regression-based models. These methods partition the range $[0,\infty)$ into consecutive non-overlapping intervals to increase the sample size per class. For instance, Xiong et al.[9] proposed DCNet, which predicts counts at multiple levels using a shared set of bins. However, this approach fails to account for the lower likelihood of large counts at lower levels, exacerbating class imbalance. Liu et al.[8] addressed this by introducing blockwise classification, where models output probability maps. However, these methods rely on Gaussian smoothing, which introduces label noise and boundary ambiguities, leading to reduced accuracy. Also, optimizing for classification accuracy alone cannot guarantee optimal performance under regression metrics such as MAE.
CLIP in Crowd Counting. CLIP [1] has demonstrated exceptional performance in recognition-related tasks, such as zero-shot image classification [1], and semantic segmentation [2], but there are only few studies in regression-related tasks, particularly crowd counting. Liang et al.[3] proposed CrowdCLIP, an unsupervised (or weakly-supervised) fully CLIP-based model, which operates on global image features extracted by CLIP’s image encoder. This global representation lacks spatial granularity and cannot capture the fine details or density variations across different parts of the image. Jiang et al.[4] extended CLIP for text-guided zero-shot object counting, incorporating additional modules such as multimodal interaction layers and decoders to generate density maps. In contrast, we propose CLIP-EBC, the first crowd counting framework fully based on CLIP, designed to estimate crowd density maps without relying on external modules. Fig. 1 presents the comparison with these methods .
III Methods
@@
Our Enhanced Blockwise Classification (EBC) framework improves the Standard Blockwise Classification (SBC) [8] from three key aspects: the quantization strategy, noise control, and loss formulation. In this section, we first briefly recapitulate SBC, and then provide a detailed explanation of our enhancements. Finally, we demonstrate how to leverage our model-agnostic EBC framework with CLIP for density map estimation, introducing our model CLIP-EBC. Fig. 2 illustrates the overview of EBC and CLIP-EBC.
III-A Enhanced Blockwise Classification
III-A1 Recap: Standard Blockwise Classification (SBC)
@@
Similar to many regression-based methods [5, 8, 6, 7], SBC also employs blockwise prediction. Formally, given an input image $\boldsymbol{X}\in\mathbb{R}_{+}^{3\times H\times W}$, with $3$ channels, height $H$ and width $W$, SBC outputs a probability map
@@
where ///// / represents the floor division operator and the integer $r$ is a model-related reduction factor. At each spatial location $(i,\,j)$, the $n$-dimensional vector $\boldsymbol{P}^{*}_{:,\,i,\,j}$ represents the predicted probability distribution of the count value in the block $(r(i-1):ri,\,r(j-1):rj)$ of the image over the $n$ predefined bins $\{\mathcal{B}_{i}\mid i=1,\cdots,n\}$. These bins satisfy
$\forall p\neq q,\,\mathcal{B}_{p}\cap\mathcal{B}_{q}=\emptyset\quad\text{and}\quad\mathcal{S}\subseteq\cup_{i=1}^{n}\mathcal{B}_{i},$
@@
III-A2 Enhancement 1: Integer-Valued Bins
@@
(a) Region with noisy labels
(b) Zoomed-in view of the region
Figure 3: An example from ShanghaiTech A dataset [12] illustrating an extremely dense area with erroneous annotations. The $8\times 8$ magenta box in Fig. 3(a) highlights a congested region with an labeled count of 9 individuals. However, the zoomed-in view in Fig. 3(b) reveals that no discernible human figures can be identified within the marked box, evidencing label noise in highly dense regions.
Integer-Valued Bins
Noise Reduction
Combined Loss
MAE $\downarrow$
140.6
88.3
85.8
✓($\lambda=0.01$) 83.9
✓($\lambda=0.10$) 83.4
✓($\lambda=1.00$) 77.9
✓($\lambda=2.00$) 82.3
TABLE I: Influence of each enhancement in our EBC framework evaluated on UCF-QNRF [10]. When no enhancement is utilized, the model deteriorates into SBC [8]. The reduced mean absolute errors demonstrate the effectiveness of our enhancement strategies and the superiority of EBC.
III-A3 Enhancement 2: Noise Reduction
@@
III-A4 Enhancement 3: Combined Loss.
Previous blockwise classification methods [8, 9] primarily focus on minimizing the cross-entropy loss between the predicted probability map $\boldsymbol{P}^{*}$ and the ground truth $\boldsymbol{P}$ but often overlook the mismatch in count values. Since two different probability distributions can yield the same classification error while having different expectations (or modes) for the predicted count values, models trained with such loss functions may not guarantee accurate count estimates. To address this limitation, we propose the Distance-Aware-Cross-Entropy (DACE) loss:
@@
where $\boldsymbol{1}$ is the indicator function, $\boldsymbol{P}$ is the one-hot encoded ground-truth probability map, $\boldsymbol{P}^{*}$ is the predicted probability map, $\boldsymbol{Y}$ is the ground-truth density map, $\boldsymbol{Y}^{*}$ is the predicted density map, and $\lambda$ is a weighting factor balancing $\mathcal{L}_{\text{class}}$ and $\mathcal{L}_{\text{count}}$. Since the ground-truth density maps are not smoothed, we use the DMCount [6] based on the discrete Wasserstein distance as the count loss $\mathcal{L}_{\text{count}}$.
@@
TABLE II: Our EBC framework can improve both classifica-tion-based and regression-based methods. We reimplement both SBC models [8] and two regression-based methods (CSRNet [5] and DMCount [6]) using our EBC framework on the UCF-QNRF dataset. Regardless of the choice of the backbone, our EBC always improves the performance.
III-B The Structure of CLIP-EBC
As CLIP-EBC is implemented under our proposed EBC framework, here we focus on how to generate the predicted probability map $\boldsymbol{P}^{*}$. Additional essential details, such as the loss function used in training, are provided in Section III-A.
where $d$ is the dimension of the CLIP embedding space.
Methods Type ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
CrowdCLIP [3] U 146.1 236.3 69.3 85.8 283.3 488.7 - -
CLIP-Count [4] R 192.6 308.4 45.7 77.4 - - - -
DMCount [6] R 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
P2PNet [13] R 52.7 85.0 6.3 9.9 85.3 154.5 77.4 362.0
CLTR [14] R 56.9 95.2 6.5 10.2 85.8 141.3 61.9 246.3
STEERER [15] R 54.5 86.9 5.8 8.5 74.3 128.3 - -
S-DCNet [9] C 58.3 95.0 6.7 10.7 104.4 176.1 - -
UEPNet [16] C 54.6 91.2 6.4 10.9 81.1 131.6 - -
CLIP-EBC (ResNet50, ours) C 54.0 83.2 6.0 10.1 80.5 136.6 38.6 90.3
CLIP-EBC (ViT/B-16, ours) C 52.5 85.9 6.6 10.5 80.3 139.3 36.6 81.7
TABLE III: Comparison of our model CLIP-EBC with recent crowd counting approaches on ShanghaiTech A & B [12], UCF-QNRF [10] and NWPU-Crowd (Val) [11]. The “Type” column groups all methods into unsupervised (“U”), regression-based (“R”), or classification-based (“C”). The best results are highlighted in bold, and the second best results are underlined. Our model CLIP-EBC demonstrates competitive performance across all datasets, with particularly significant superiority on the NWPU validation split, where it outperforms all existing methods. : Results of cross-dataset evaluation.